OpenStudy (anonymous):

what is the square root of 4*36 +5

OpenStudy (anonymous):

\[ M = 36 \sqrt{4m}+5\]

OpenStudy (anonymous):

I have to evaluate the that question.

OpenStudy (anonymous):

@halorazor

OpenStudy (anonymous):

You tagged it wrong xD but I got you.

OpenStudy (anonymous):

Spelled it wrong but it's okay. ;p I spell it weird.

OpenStudy (anonymous):

What exactly are you trying to figure out?

OpenStudy (anonymous):

Evaluate the following expression for m = 36 \[\sqrt{4m}+5\]

OpenStudy (anonymous):

Well \[\sqrt{4(36)} +5\] \[\sqrt{144}+5\] \[12+5\] \[17\]

OpenStudy (anonymous):

Thank you! Do you mind helping me on a few more?

OpenStudy (anonymous):

I'd be more than happy to.

OpenStudy (anonymous):

And yeah, it was a bit confusing at first Peach xD

OpenStudy (anonymous):

I have to simplify. \[\sqrt{6} (6+\sqrt{3})\]

OpenStudy (anonymous):

I mean I think you would distribute it out.

OpenStudy (anonymous):

Well how would I do that?

OpenStudy (anonymous):

Multiply the Square root of 6 to the 6 and the square root of 3.

OpenStudy (anonymous):

I don't think I'm right bur is it 24.6?

OpenStudy (anonymous):

It said simplify. :p not calculate.

OpenStudy (anonymous):

I'm confused.